OpenStudy (anonymous):

Find the value of 10 sqrt 2 / sqrt 5 *i picked 2sqrt 5?

OpenStudy (anonymous):

\[10 * \sqrt{2} * \sqrt{5} /\sqrt{5} * \sqrt{5}\] \[10 * \sqrt{2 *5}/ \sqrt{5 * 5}\] \[10 * \sqrt{10}/ \sqrt{25}\] (sqrt of 25 is 5) \[10 * \sqrt{10}/5\] 5 cancels out because 2 * 5 is 10 \[2 * \sqrt{10}\] Hope that helps! n_n